OpenStudy (anonymous):

\[\sqrt{30 - x} = x\] :)

OpenStudy (anonymous):

\[\sqrt{30 - x}^{2}\] = x ^2 30 - x = x^2 x^2 + x - 30 = 0

OpenStudy (anonymous):

then factorize it ..